### v1

# Bushfires and Balance: Proactive versus Reactive Policies in Prescribed Burning

Abstract.
We introduce a model to describe the interplay between prescribed burning and bushfires based on a system of ordinary differential equations.
We show that the system possesses a unique steady-state, whose stability depends on the policy governing prescribed burning. Specifically, a Reactive Policy in which prescribed burning activities are increased in response to bushfire events risks to produce instability and requires an additional feedback control, while a Proactive Policy in which moderate prescribed burning replaces natural bushfires may be more effective to maintain a balanced state. A high regeneration factor is also helpful to ensuring the stability of the equilibrium state.
The results presented here should not be regarded as a definitive solution to the issue of prescribed burning, as numerous specific factors must be considered in any comprehensive policymaking approach. Nonetheless, the stability analysis developed in this paper suggests that policies driven primarily by reactive responses to bushfire events may be emotionally charged and prone to instability. These reactive policies would benefit from incorporating feedback control mechanisms. In contrast, strategies based on proactive prescribed burning in moderate quantities, focused on prevention rather than reaction, are likely to be more effective in maintaining long-term ecological stability.
Key words and phrases:
bushfire models; controlled burning; stability analysis; proactive management.
2020 Mathematics Subject Classification:
34A34; 34D05; 34D20; 34H15; 92-10.
Supported by Australian Research Council DP250101080 and DE190100379.
1. Introduction
Prescribed burns are a key strategy for managing bushfire risks by reducing fuel loads. The relationship between the frequency of bushfires and the implementation of prescribed burns is influenced by various factors, such as environmental conditions and resource availability. Moreover, as detailed for instance in Section 1.3 of [REPO], “prescribed burning can be a contentious issue given that management objectives for its use often conflict with other land management objectives (e.g. effective fuel management can alter habitat features), and its use can have a range of off-site consequences (e.g. smoke can cause discomfort and disruption to affected communities)”, and opposition from communities can rise “based on a view that prescribed burning is environmentally destructive, unnecessary, and probably ineffective in reducing bushfire threat”. Adverse weather conditions, such as extreme heat and dryness, may also limit the windows for safe prescribed burning, and extreme conditions could lead to dangerous fire seasons “regardless of treatment rate” (see e.g. [CLARKE]).
The literature about controlled burning has a very long tradition (see e.g. [q-iw0pf294rf8vj30276jgrth]) and prescribed burning was trialled as early as the 1890s for forest regeneration purposes (see [iadjhfw0ireuy03245-63FSp12-35oyu]). Notably, indigenous peoples worldwide have employed burning practices long before modern scientific documentation (see e.g. [BUTZ09, qa0DOFHWP3REP9H203EPFL9yusdhfiovlf30er-21, amo22]). In general, the topic of hazard reduction burns is quite controversial and is often the subject of intense, public debate. Critics consider it the byproduct of short-term thinking, whose “benefits are greater in the immediate aftermath”, but ineffective in the presence of adverse conditions and potentially dangerous since they may “activate seed banks when they burn, which in turn leads to the growth of more flammable vegetation below”, see [ojdsacnpreoogub987381i2kwertigf]. The impact of controlled burning on air quality and biodiversity is also a subject of environmental concern (see e.g. [Haikerwal04052015, BIOD]).
Given the complexity of the problem and its practical importance, it is desirable to have a quantitative model assessing the relationship between bushfire intensity and preventive burns. In terms of available data, determining the variations of prescribed burns following major bushfire events is challenging. This is because of variations in reporting and implementation across different regions and time periods. As a rule of thumb, however, it seems that recommendations often call for increased prescribed burning after major bushfires. This is especially true after major media coverage. For example, after the 2009 Black Saturday Bushfires the Victorian Bushfires Royal Commission recommended a doubling of the prescribed burning program, increasing preventive burning to 5% of public land annually (see Recommendation 56 in [RECOMM]) and this advice was broadly reported in the media (see e.g. [m2edia]). Along the same lines, in the United States, a significant rise in controlled burns occurred as a response to massive fires in California, Oregon, and other Western states, such as the catastrophic wildfires in 2017 and 2018 (like the Camp Fire, which destroyed Paradise, CA). As a response to these events, in 2020 approximately 9.4 million acres were treated with prescribed fire in the United States. The Forest Service aims to treat up to an additional 20 million acres across the National Forest System over the next decade (see [2021REP]).
The simple model that we present here accounts for the evolution of three positive quantities: the amount of burning regions in a given area due to bushfire, the available amount of land which can be potentially set on fire, and the portion of the area which is burnt intentionally to prevent bushfire.
These quantities are functions of time (though, in a sense, they have to be thought as “averaged” quantities over some interval of time, referring to a given region of reference, since the objective here is to understand the effectiveness of a bushfire prevention plan, which typically ranges over a sufficiently long time elapse).
We will show that the model possesses a unique fixed point. The stability of this equilibrium depends on the policy adopted for the prescribed burning. Specifically, we will show that when prescribed burning activities are incremented as a result of bushfire events, the equilibrium is unstable. Conversely, when the (moderate) prescribed burning replaces natural bushfires, the equilibrium is stable.
2. Mathematical setting and main results
2.1. Modeling bushfires and prescribed burning
In our notation, $a$ denotes the amount of unburnt land available at a given time $t$, $b$ represents the amount of land subject to preventive controlled burns, and $f$ accounts for the frequency and intensity of bushfires.
We suppose that the dynamics of this system is regulated by the following system of three ordinary differential equations. The first equation encodes the change in available land, which decreases due to bushfires and controlled burns and increases thanks to a regeneration factor. Specifically, we suppose that the available land decreases proportionally with respect to controlled burns, while the bushfires effect is linear with respect to both the available land and the bushfire intensity. As for the regeneration effect, we assume that it is proportional to the amount of burnt land (i.e., normalising to $1$ the total amount of land, to $1-a$). These considerations lead to the equation
$\dot{a}=-\alpha fa-\beta b+\gamma(1-a),$
where $\alpha$, $\beta$, $\gamma\in(0,+\infty)$.
The second structural equation that we consider describes the change of bushfire frequency and intensity due to the availability of fuel (assumed to be proportional to the available land), allowing for a natural decay over time, namely
(2.1) $\dot{f}=\zeta a-\eta f,$
where $\zeta$, $\eta\in(0,+\infty)$.
The third equation models the variation of controlled burns. For this, we suppose that there exists a suitable bushfire intensity threshold $f_{0}\in(0,+\infty)$ which represents the “optimal” amount of fires needed in nature for beneficial purposes (such as regenerating the landscape, removing invasive weeds, releasing nutrients from dead plants and animals into the soil, creating habitats such as hollows in trees and logs, and cracking seed coats to trigger germination).
Thus, we suppose that the change in controlled burns is proportional to the available land, as well as to the excessive bushfire intensity. These observations lead to the equation
(2.2) $\dot{b}=\theta(f-f_{0})a,$
where $\theta\in\mathbb{R}\setminus\{0\}$.
Summarising, the model comprises the system of ordinary differential equations
(2.3) $\begin{cases}&\dot{a}=-\alpha fa-\beta b+\gamma(1-a),\\ &\dot{f}=\zeta a-\eta f,\\ &\dot{b}=\theta(f-f_{0})a.\end{cases}$
In spite of its simplicity, this system of equations requires a tailored analysis, since it does not fall into standard categories (for example, no choice of parameters reduces it to a gradient system).
It is readily seen that the only physically relevant (i.e., corresponding to positive values of the available land $a$) equilibrium for this system is
(2.4) $a_{\star}:=\frac{\eta f_{0}}{\zeta},\quad f_{\star}:=f_{0}\quad{\mbox{and}}\quad b_{\star}:=\frac{\gamma\zeta-\eta f_{0}(\gamma+\alpha f_{0})}{\beta\zeta}.$
This equilibrium is already of practical interest. The relation between $a_{\star}$ and $f_{\star}$, that is,  $f_{\star}=\frac{\zeta a_{\star}}{\eta}$, quantifies the optimal bushfire intensity threshold in terms of the environmental land. The third relation in (2.4), being equivalent to $b_{\star}=\gamma-\frac{a_{\star}(\gamma\eta+\alpha\zeta a_{\star})}{\beta\eta}$, determines the optimal preventive burning given the environmental land.
For consistency, one expects the available area at equilibrium $a_{\star}$ to be bounded from above by the area of the full region, which has been normalised to $1$. This implies
(2.5) $\eta f_{0}\leqslant\zeta.$
On a similar note, one expects the burning area to be nonnegative and less than or equal to the full available region, namely $b_{\star}\in[0,a_{\star}]$, which entails
(2.6) $(\gamma+\alpha f_{0})\eta f_{0}\leqslant\gamma\zeta\leqslant(\beta+\gamma+\alpha f_{0})\eta f_{0}.$
The uniqueness of the equilibrium in (2.4) identifies the steady state ensuring safe management of the land. However, to make this equilibrium “robust”, it would be desirable that small deviations do not make the system drift away from the steady state. That is, we require a stability analysis of the equilibrium, in which the sign of the parameter $\theta$ will play a crucial role – see Theorems 2.1 and 2.3.
Regarding the units of measurement in our model, in addition to traditional land area metrics (e.g., acres), we mention that wildfire frequency can be expressed in acres per year (total area affected over time), years (average fire return interval at a location), or fires per unit area per year (a proxy for regional fire frequency). Similarly, wildfire intensity can be measured in megawatts, representing the total power emitted as thermal radiation, often assessed via satellite data. However, we emphasise that the key advantage of our model is its applicability as a conceptual framework, based on the overall structure of the problem more than on specific details, providing policymakers with early warnings about potential structural instabilities in policies set in place without a long-term plan.
2.2. Reactive Policy: Increasing prescribed burning after bushfires
A natural assumption is that the parameter $\theta$ in (2.2) is positive. The ansatz in this case is that the more severe the bushfire events, the more extensive the preventive activity, as an outcome of committee recommendations, specific policies, increased public sensibility, and media coverage. Due to this, we consider the case $\theta>0$ corresponding to Reactive Policy.
Figure 1. Number of prescribed burns and wildfires in different areas of Western Australia, 2019–2022. Source: Brodie James Loller, 2022, %****␣unstable250221.tex␣Line␣325␣****https://storymaps.arcgis.com/stories/6428d63b54e4406eadb62e633f5f41fb.
Interestingly, under this condition, the equilibrium in (2.4) is linearly unstable:
Theorem 2.1.
The equilibrium in (2.4) is linearly unstable when $\theta\in(0,+\infty)$, since the Jacobian matrix associated with the system in (2.3) and evaluated at (2.4) always possesses at least one real, positive eigenvalue.
More precisely, depending on the structural parameters $\alpha$, $\beta$, $\gamma$, $\zeta$, $\eta$, $\theta$, and $f_{0}$, this matrix can present:
  • three real eigenvalues, of which one is positive and two are negative; or
  • one real, positive eigenvalue and two distinct, complex conjugate eigenvalues with negative real part.
There exist parameters that realise each of the cases above.
We observe that, as a byproduct of the eigenvalue classification in Theorem 2.1, it follows that the unstable space of the equilibrium in (2.4) has dimension one (namely, it is spanned by the only positive eigenvector).
For practical purposes, the statement of Theorem 2.1, highlighting the instability of the only physical equilibrium of the system, could be, in principle, interpreted as a negative result, hinting at significant concrete difficulties in managing bushfires merely on the basis of preventive burns. However, the quantitative analysis and the eigenvalue characterisation in Theorem 2.1 allows one to efficiently utilise an “augmentation method” or “dynamic extension”: more specifically, one can introduce an additional ordinary differential equation to the system to stabilise the equilibrium. This additional feedback mechanism influences the dynamics and provides a stabilising damping effect:
Theorem 2.2.
There exists a feedback mechanism which, added to (2.3), makes the equilibrium in (2.4) linearly stable.
While the statement proposed in Theorem 2.2 may sound vague and non-specific, the structure provided in Theorem 2.1 actually permits an explicit construction of this feedback mechanism, based on a partial diagonalisation of the Jacobian matrix and the coupling with a suitable, linear reaction term: in this sense, the proof of Theorem 2.2 is more interesting than its statement, since it exhibits a “constructive” procedure to stabilise the fixed point. Indeed, from a mathematical perspective, this feedback mechanism penalises the positive eigenvalue direction, making its identification useful for wildfire management. Furthermore, recognising the potential risk associated with instability can help justify additional preventive measures to regulate fuel availability.
Figure 2. Trends in (a) prescribed fire area (hectares burned), (b) wildfire incidence (number of unplanned ignitions) and (c) wildfire area (hectares burned by wildfire) from 1982 to 2012 at Fort Benning, GA. Source: Figure 1 in [iwfkhvgri0eotAsd].
2.3. Proactive Policy: Increasing (with moderation) prescribed burning in the absence of bushfires
Interestingly, another method to stabilise the equilibrium in (2.4) consists in modifying the strategy of the preventive burning. For example, one could perform less controlled burning in the presence of a highly intense natural fire activity. The ansatz in this case is that the severe bushfire events have already removed the available fuel from the soil, hence less preventive activity is necessary. Low bushfire activity leaves excessive fuel in the land which needs to be preventively removed via controlled burning.
Bushfires may be the outcome of extreme weather conditions, making controlled burns unsafe or impossible, which may also suggest that a proactive policy needs to decrease prescribed burning when bushfires are active.
This controlled burning strategy would maintain the form of (2.2), changing only the sign of its right-hand-side, leading to
$\dot{b}=-\vartheta(f-f_{0})a,$
with $\vartheta:=-\theta\in(0,+\infty)$.
From the mathematical standpoint, this plan leaves the system of equations in (2.3) formally unchanged, except that now $\theta=-\vartheta$ is negative (we stress that the equilibrium in (2.4) and conditions (2.5) and (2.6) do not involve the parameter $\theta$ and hence remain unaffected by its sign).
Figure 3. The evolution of forest fires and prescribed burning in the Massif des Madres. Source: Figure 9 in [LAMBE].
The relevant fact is that, with this program of preventive fires, the equilibrium (2.4) becomes stable for suitable parameter ranges:
Theorem 2.3.
The equilibrium in (2.4) is linearly stable for the system in (2.3) when $\theta=-\vartheta\in(-\infty,0)$, as long as
(2.7) $\left(\alpha f_{0}+\gamma+\eta\right)\left(2\alpha f_{0}\eta+\gamma\eta\right)>\beta\eta\vartheta f_{0},$
since the eigenvalues associated with the Jacobian matrix each have negative real parts.
More precisely, under assumption (2.7), depending on the structural parameters $\alpha$, $\beta$, $\gamma$, $\zeta$, $\eta$, $\theta$, and $f_{0}$, this matrix can present:
  • three real, negative eigenvalues,
  • one real, negative eigenvalue and two complex conjugate eigenvalues with negative real part.
There exist parameters that realise each of the cases above.
When instead
$\left(\alpha f_{0}+\gamma+\eta\right)\left(2\alpha f_{0}\eta+\gamma\eta\right)<\beta\eta\vartheta f_{0},$
the equilibrium in (2.4) is linearly unstable for the system in (2.3) and the associated Jacobian matrix presents one real, negative eigenvalue and two complex conjugate eigenvalues with positive real part.
Finally, when
$\left(\alpha f_{0}+\gamma+\eta\right)\left(2\alpha f_{0}\eta+\gamma\eta\right)=\beta\eta\vartheta f_{0},$
the linear stability of the system is undetermined, since the Jacobian matrix possesses one real, negative eigenvalue and two complex, purely imaginary eigenvalues.
Condition (2.7) is interesting since, being satisfied for small values of $\vartheta$, it suggests that moderate (rather than excessive) controlled burning could provide more stable outcomes. We can further determine the relevant factors in making decisions on controlled burning. In environments with low levels of natural regeneration, less prescribed burning is desirable, as the factor on the left-hand side of (2.7) will be lower. If the regenerative factor $\gamma$ is high, however, or the optimal bushfire intensity $f_{0}$ is high, then condition (2.7) is satisfied and more controlled burning remains permissible without sacrificing stability.
Regarding Theorems 2.1 and 2.3, we observe that, while structural parameters are often difficult to measure in practice, a key advantage of our analysis is that it relies primarily on a single parameter, namely $\theta$, and, crucially, only on its sign. In this spirit, although distinguishing between “reactive” and “proactive” strategies is often impractical—being implicitly tied to the time scale used to differentiate “actions” from “reactions”—identifying the underlying blueprint of these ideal programs and assessing their viability through rigorous, quantitative analysis can provide valuable insights into the fundamental structure of a complex phenomenon (see Figures 1, 2, and 3 for several comparisons between natural bushfires and controlled burning activities which possibly hint at different management strategies).
2.4. Further observations
Regarding the condition $b\leqslant a$ describing the fact that the land devoted to preventive burning is less than or equal to the total available land, a natural question is whether all solutions of (2.3) fulfill this prescription, e.g. whether or not when the initial condition satisfies $b(0)<a(0)$ one has that $b(t)\leqslant a(t)$, or at least whether or not this is true under suitable assumptions on the structural parameters $\alpha$, $\beta$, $\gamma$, $\zeta$, $\eta$, $\theta$, and $f_{0}$. The next result states that this is not the case:
Theorem 2.4.
Given any $\alpha$, $\beta$, $\gamma$, $\zeta$, $\eta$, $\theta$, and $f_{0}$ as above, there exist initial conditions $a(0)\in(0,1)$, $b(0)\in\left(0,a(0)\right)$, and $f(0)\in(0,+\infty)$ and a time $T>0$ such that the solution of (2.3) satisfies
$10<a(T)<b(T)<$ 1 .
The situation put forth in Theorem 2.4 is interesting for practical applications, since it highlights the risk of exhausting the available land due to (a combined effect of bushfires and) preventive burning.
While, according to Theorem 2.4, the condition $b<a$ is not preserved under the flow of (2.3), the conditions $f>0$ and $a\in(0,1)$ do persist (as long as $0\leqslant b\leqslant a$):
Theorem 2.5.
Let $\alpha$, $\beta$, $\gamma$, $\zeta$, $\eta$, $\theta$, and $f_{0}$ be as above. Suppose that, for some $T>0$, the solution of (2.3) in the interval $[0,T]$ is such that $a(t)\in(0,1)$, $f(t)\in(0,+\infty)$, and $b(t)\in\big{[}0,a(t)]$ for all $t\in(0,T)$.
Then, $a(T)\in(0,1)$ and $f(T)\in(0,+\infty)$.
We stress that Theorems 2.4 and 2.5 are valid both when $\theta\in(0,+\infty)$ and when $\theta\in(-\infty,0)$.
2.5. Caveats and perspectives
We stress that, as is usual when dealing with the potential implications of theoretical works on concrete scenarios, scientific results must be correctly interpreted, and a pinch of salt is necessary when attempting to translate them into actual policies. On the one hand, one of the advantages of the model (2.3) lies in its intrinsic simplicity and the relatively small number of structural parameters involved in describing complex interactions. This allows for a thorough analytical treatment with explicit quantifications. On the other hand, the model (2.3) is a simplification of the highly complex, high-dimensional dynamics involved in bushfire management. For instance, the equation modeling bushfire frequency in (2.1) can be enriched by incorporating a seasonal term (this can be achieved by adding a periodic forcing term to its right-hand side). We wish to also mention two other specific factors. Latent flammability (see [LINZYL]) and biodiversity (see [BIOD]), can each be impacted by hazard reduction, leading to another feedback effect that could influence long-time dynamics.
Moreover, other key factors could be considered in the system (2.3), such as public awareness, public perception of risks, involvement of local stakeholders, political trends, media coverage, climate considerations, impact on the quality of air, alternative managing procedures such as selective logging and mulching practices, and restoration activities (which themselves evolve over time and both influence and are influenced by bushfire events). To study the combination of these factors, it will also be important to distinguish the different time scales at which they take place (for example, one can model both regular and catastrophic wildfire events as either random or periodic, but their frequency differs, and climate change is causing the period to shrink and the frequency to increase). Although incorporating all these factors would likely hinder a fully analytical treatment of the problem, we plan to revisit these considerations in a future work.
2.6. Organisation of the paper
The rest of this paper is organised as follows. The proofs of the above results are presented in Section 3, with the demonstration of some auxiliary algebraic results deferred to the appendix. Some conclusions will be summarised at the end of the paper.
3. Proofs of the main results
3.1. Some preliminary algebraic observations
In the proof of Theorems 2.1 and 2.3, when establishing that the system may possess, depending on the paramenters, real eigenvalues as well as complex ones, the following observations will come in handy:
Lemma 3.1.
Let $B$, $C$, $D\in(0,+\infty)$, with $BC\leqslant 3D$. Then, the polynomial $\lambda^{3}+B\lambda^{2}+C\lambda+D$ possesses two complex conjugate roots.
If also $BC<D$, then the real part of these roots is positive.
Lemma 3.2.
Let
(3.1) $B:=\alpha f_{0}+\gamma+\eta,\qquad C:=2\alpha f_{0}\eta+\gamma\eta\qquad{\mbox{and}}\qquad E:=-\beta\eta\theta f_{0}.$
Let also
(3.2) $\triangle:=B^{2}C^{2}-4C^{3}-4B^{3}E-27E^{2}+18BCE.$
Given $\theta\in\mathbb{R}\setminus\{0\}$, there are suitable choices of the structural parameters $\alpha$, $\beta$, $\gamma$, $\zeta$, $\eta$, and $f_{0}$, fulfilling conditions (2.5), (2.6) and $BC>E$, which reads
(3.3) $\left(\alpha f_{0}+\gamma+\eta\right)\left(2\alpha f_{0}\eta+\gamma\eta\right)>-\beta\eta\theta f_{0},$
allowing for both of the possibilities $\triangle>0$ and $\triangle<0$.
For the convenience of the reader, the proof of these results is deferred to the appendix. Notice that the parameter $\theta$ does not play a role in Lemma 3.2, allowing us to use the above result both when $\theta\in(0,+\infty)$ and when $\theta\in(-\infty,0)$.
In our framework the interest of Lemma 3.2 is that it detects the sign of the discriminant $\triangle$ of the cubic equation
$\lambda^{3}+B\lambda^{2}+C\lambda+E=0,$
where $B$, $C$ and $E$ are as in (3.1).
Specifically, if the discriminant is positive, the cubic equation has three distinct real roots, while if the discriminant is negative, the cubic equation has one real root and two complex conjugate roots (see e.g. Section 6.2 in [MR1409810] for the role of discriminants in polynomial equations). Therefore, Lemma 3.2 will allow us to conclude that both the case of real and complex eigenvalues occur in our system, according to the statements of Theorems 2.1 and 2.3.
3.2. Proof of Theorem 2.1
As is customary, the stability of an equilibrium is determined by the Jacobian matrix (at the equilibrium). In this setting, one says that the equilibrium is linearly stable if all eigenvalues have negative real parts (this, in fact, ensures that the fixed point is locally attractive with exponential strength, see e.g. Proposition 6 on page 378 of [MR698947]) and is linearly unstable if at least one of the eigenvalues has a positive real part.
The Jacobian matrix associated with the system of ordinary differential equations in (2.3) takes the form
(3.4) $J(a,f,b):=\begin{bmatrix}-\alpha f-\gamma&-\alpha a&-\beta\\ \zeta&-\eta&0\\ \theta(f-f_{0})&\theta a&0\end{bmatrix}.$
This and (2.4) give that
(3.5) $J(a_{\star},f_{\star},b_{\star})=\begin{bmatrix}-\alpha f_{0}-\gamma&-\frac{\alpha\eta f_{0}}{\zeta}&-\beta\\ \zeta&-\eta&0\\ 0&\frac{\eta\theta f_{0}}{\zeta}&0\end{bmatrix}.$
The eigenvalues of this matrix are determined by the third-degree characteristic equation
(3.6) $P(\lambda):=\lambda^{3}+\left(\alpha f_{0}+\gamma+\eta\right)\lambda^{2}+\left(2\alpha f_{0}\eta+\gamma\eta\right)\lambda-\beta\eta\theta f_{0}=0.$
Key for our analysis is that the characteristic polynomial has the form
(3.7) $\begin{split}&P(\lambda)=\lambda^{3}+B\lambda^{2}+C\lambda-D,\\ &{\mbox{with\leavevmode\nobreak\ $B$, $D\in(0,+\infty)$ and\leavevmode\nobreak\ $C\in\mathbb{R}$.}}\end{split}$
In this case we know that $C\in(0,+\infty)$, but we are not using this information in what follows.
Since $P(0)=-D<0$ and $P(+\infty)=+\infty$, there exists at least one root $\lambda_{1}$ of $P$ lying in the interval $(0,+\infty)$, which proves that the equilibrium in (2.4) is linearly unstable.
Let us now consider the other (possibly complex) roots of $P$, denoted by $\lambda_{2}$ and $\lambda_{3}$. Since
$\lambda^{3}+B\lambda^{2}+C\lambda-D=P(\lambda)=(\lambda-\lambda_{1})(\lambda-\lambda_{2})(\lambda-\lambda_{3})$
$\qquad=\lambda^{3}-(\lambda_{1}+\lambda_{2}+\lambda_{3})\lambda^{2}+(\lambda_{1}\lambda_{2}+\lambda_{1}\lambda_{3}+\lambda_{2}\lambda_{3})\lambda-\lambda_{1}\lambda_{2}\lambda_{3},$
it follows that
$-B=\lambda_{1}+\lambda_{2}+\lambda_{3}\qquad{\mbox{and}}\qquad D=\lambda_{1}\lambda_{2}\lambda_{3}.$
Since $\lambda_{1}\in(0,+\infty)$, we deduce that
(3.8) $\lambda_{2}\lambda_{3}=\frac{D}{\lambda_{1}}\in(0,+\infty)$
and that
(3.9) $\lambda_{2}+\lambda_{3}=-B-\lambda_{1}\in(-\infty,0).$
Now we distinguish two cases: either $\lambda_{2}$, $\lambda_{3}$ are real, or are complex conjugate. If they are real, it follows from (3.9) that at least one of them is negative, say $\lambda_{2}\in(-\infty,0)$. Hence, by (3.8), also $\lambda_{3}=\frac{D}{\lambda_{1}\lambda_{2}}\in(-\infty,0)$.
If instead $\lambda_{2}$, $\lambda_{3}$ are complex conjugate, we write $\lambda_{2}=x+iy$ and $\lambda_{3}=x-iy$, for some $x\in\mathbb{R}$ and $y\in\mathbb{R}\setminus\{0\}$, and we deduce from (3.9) that $x\in(-\infty,0)$.
This completes the eigenvalue analysis stated in Theorem 2.1.
Lemma 3.2 guarantees that both of the above scenarios are possible, depending on the choice of the parameters. This finishes the proof of Theorem 2.1
3.3. Proof of Theorem 2.2
We denote by $y$ the (column) vector with entries $a$, $f$, $b$ (the corresponding equilibrium in (2.4) will be denoted by $y_{\star}$). Hence, we rephrase (2.3) in the form
(3.10) $\dot{y}=Y(y),$
for a suitable (nonlinear) map $Y:\mathbb{R}^{3}\to\mathbb{R}^{3}$. In this setting, the equilibrium property of $y_{\star}$ reads $Y(y_{\star})=0$.
We use the setting of (3.5) with the short notation $J_{\star}:=J(a_{\star},f_{\star},b_{\star})$. By virtue of Theorem 2.1, we consider the eigenvalues $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ of $J_{\star}$, with $\lambda_{1}\in(0,+\infty)$ and $\Re\lambda_{2}$, $\Re\lambda_{3}\in(-\infty,0)$.
We perform a Schur triangulation (see e.g. Section 2.3 in [MR2978290]), namely we take a unitary matrix $Q$ such that
$J_{\star}=Q^{-1}UQ,$
where $U$ is an upper triangular matrix of the form
$U=\left(\begin{matrix}\lambda_{1}&*&*\\ 0&\lambda_{2}&*\\ 0&0&\lambda_{3}\end{matrix}\right).$
Here, “*” denotes some complex entry (possibly different from one another).
We consider the change of variable $x:=Qy$ and define
$x_{\star}:=Qy_{\star}\qquad{\mbox{and}}\qquad X(x):=QY(Q^{-1}x).$
In this way, equation (3.10) becomes
(3.11) $\dot{x}=Q\dot{y}=QY(y)=QY(Q^{-1}x)=X(x),$
and note that $X(x_{\star})=QY(y_{\star})=0$.
When needed, we will use coordinates by writing $x_{j}$ for the $j$th entry of the vector $x$ and $X_{j}$ for the $j$th entry of the vector $X$.
In this scenario, we have that $x_{\star}$ is an equilibrium for the system of ordinary differential equations in (3.11), which is in turn equivalent to the original system in (2.3) up to a linear mapping. The corresponding Jacobian matrix satisfies
(3.12) $\begin{split}&\frac{\partial X}{\partial x}(x_{\star})=Q\frac{\partial Y}{\partial y}(Q^{-1}x_{\star})\,Q^{-1}=Q\frac{\partial Y}{\partial y}(y_{\star})\,Q^{-1}\\ &\qquad=QJ_{\star}\,Q^{-1}=U=\left(\begin{matrix}\lambda_{1}&*&*\\ 0&\lambda_{2}&*\\ 0&0&\lambda_{3}\end{matrix}\right).\end{split}$
To stabilise the equilibrium $x_{\star}$ (or equivalently, up to a linear mapping, to stabilise the original equilibrium $y_{\star}$), we propose the following system of ordinary differential equations which includes now a feedback function $\omega:\mathbb{R}\to\mathbb{R}$:
(3.13) $\begin{cases}\dot{x}_{1}=X_{1}\big{(}x_{1}-\tau(\omega-x_{\star,1}),x_{2},x_{3}\big{)},\\ \dot{x}_{2}=X_{2}(x),\\ \dot{x}_{3}=X_{3}(x),\\ \dot{\omega}=\sigma(x_{1}-\omega).\end{cases}$
Above $\sigma$, $\tau\geqslant 0$ represent feedback parameters. When $\tau=0$, the system (3.13) is equivalent to the original one. When $\sigma$, $\tau>0$, the feedback function $\omega$ evolves, accounting for the unstable direction $x_{1}$ and acts on its evolution (the stable directions $x_{2}$ and $x_{3}$ are not affected by the feedback).
We stress that $x_{\star\star}:=(x_{\star},x_{\star,1})$ is an equilibrium for (3.13) and it follows from (3.12) that the corresponding Jacobian matrix has the form
$J_{\star\star}:=\left(\begin{matrix}\lambda_{1}&*&*&-\tau\lambda_{1}\\ 0&\lambda_{2}&*&0\\ 0&0&\lambda_{3}&0\\ \sigma&0&0&-\sigma\end{matrix}\right).$
The corresponding eigenvalues are therefore
$\lambda_{2},\quad\lambda_{3},\quad\frac{\lambda_{1}-\sigma+\sqrt{(\sigma+\lambda_{1})^{2}-4\sigma\tau\lambda_{1}}}{2},\quad{\mbox{and}}\quad\frac{\lambda_{1}-\sigma-\sqrt{(\sigma+\lambda_{1})^{2}-4\sigma\tau\lambda_{1}}}{2}.$
The real part of all these eigenvalues is negative as long as $\sigma>\lambda_{1}$ and $\tau>\frac{(\sigma+\lambda_{1})^{2}}{4\sigma\lambda_{1}}$, thus providing the desired stability property. The proof of Theorem 2.2 is thereby complete.
3.4. Proof of Theorem 2.3
The matrix analysis in Section 3.2 can be used here with $\theta=-\vartheta\in(-\infty,0)$ and accordingly, by (3.6), the characteristic polynomial takes the form
$P(\lambda):=\lambda^{3}+\left(\alpha f_{0}+\gamma+\eta\right)\lambda^{2}+\left(2\alpha f_{0}\eta+\gamma\eta\right)\lambda+\beta\eta\vartheta f_{0}.$
In particular,
$\begin{split}&P(\lambda)=\lambda^{3}+B\lambda^{2}+C\lambda+D,\\ &{\mbox{with\leavevmode\nobreak\ $B$, $C$, $D\in(0,+\infty)$.}}\end{split}$
This is a cubic polynomial and thus it possesses at least one real root $\lambda_{1}$. Since all the coefficients of $P$ are positive,
(3.14) all real roots must be negative.
Now, we distinguish three cases: either
(3.15) $\left(\alpha f_{0}+\gamma+\eta\right)\left(2\alpha f_{0}\eta+\gamma\eta\right)>\beta\eta\vartheta f_{0},$
or
(3.16) $\left(\alpha f_{0}+\gamma+\eta\right)\left(2\alpha f_{0}\eta+\gamma\eta\right)<\beta\eta\vartheta f_{0},$
or
(3.17) $\left(\alpha f_{0}+\gamma+\eta\right)\left(2\alpha f_{0}\eta+\gamma\eta\right)=\beta\eta\vartheta f_{0}.$
We show that
(3.18) when (3.15) holds true, the equilibrium is linearly stable.
To this end, in view of (3.14), it suffices to consider the case in which $P$ presents two complex conjugate roots $\lambda_{2}=x+iy$ and $\lambda_{3}=x-iy$, with $x\in\mathbb{R}$ and $y\in\mathbb{R}\setminus\{0\}$.
We can rewrite (3.15) as $BC>D$ and therefore
$P(-B)=-B^{3}+B^{3}-BC+D<0.$
Since $P(+\infty)=+\infty$, this gives that
(3.19) $\lambda_{1}\in(-B,0).$
Now, since
(3.20) $\begin{split}&\lambda^{3}+B\lambda^{2}+C\lambda+D=P(\lambda)=(\lambda-\lambda_{1})(\lambda-\lambda_{2})(\lambda-\lambda_{3})\\ &\qquad=\lambda^{3}-(\lambda_{1}+\lambda_{2}+\lambda_{3})\lambda^{2}+(\lambda_{1}\lambda_{2}+\lambda_{1}\lambda_{3}+\lambda_{2}\lambda_{3})\lambda-\lambda_{1}\lambda_{2}\lambda_{3}\end{split}$
we thus have
(3.21) $2x=\lambda_{2}+\lambda_{3}=\lambda_{1}+\lambda_{2}+\lambda_{3}-\lambda_{1}=-B-\lambda_{1}.$
It follows from this and (3.19) that $x<0$, which completes the proof of (3.18).
Further, by virtue of Lemma 3.2, we have that both the case of three real roots and that of one real root and two complex conjugate roots can occur. This completes the case in which (3.15) is satisfied.
Let us now consider the case in which (3.16) holds true. In this situation we have that $BC<D$ and thus, by (3.14) and Lemma 3.1, we know that the characteristic polynomial presents a real, negative root and two complex conjugate roots with positive real part, also giving that the equilibrium is linearly unstable.
Finally, if (3.17) holds true, the linear stability of the system is undetermined, since in this case $BC=D$ and thus
$P(\lambda)=\lambda^{3}+B\lambda^{2}+C\lambda+BC=(\lambda+B)(\lambda^{2}+C),$
which presents two purely imaginary roots. This finishes the proof of Theorem 2.3
3.5. Proof of Theorem 2.4
Consider the initial data $a(0):=b(0)\in\left(\frac{\gamma}{\alpha f_{0}+\beta+\gamma},1\right)$ and $f(0):=f_{0}$. Then, we deduce from (2.3) that
$\dot{b}(0)-\dot{a}(0)=\theta(f(0)-f_{0})a(0)+\alpha f(0)a(0)+\beta b(0)-\gamma(1-a(0))$
$\quad=\big{(}\alpha f(0)+\beta+\gamma\big{)}a(0)-\gamma>0.$
As a result, there exists $\varepsilon>0$ small enough such that $b(\varepsilon)-a(\varepsilon)>0>b(-\varepsilon)-a(-\varepsilon)$, with $a(\varepsilon)$, $b(\varepsilon)\in(0,1)$. The desired result now follows via a time translation $t\mapsto t+\varepsilon$, by choosing $T:=2\varepsilon$
3.6. Proof of Theorem 2.5
We argue for the sake of contradiction. Three cases can occur: either $a(T)=0$, $a(T)=1$, or $f(T)=0$.
If $a(T)=0$ for some $T>0$, then $b(T)\leqslant a(T)=0$ and thus $\dot{a}(T)=-\beta b(T)+\gamma>0$. But then, for $\varepsilon>0$ small enough, $a(T-\varepsilon)<a(T)=0$, in contradiction with our assumptions.
Similarly, if $f(T)=0$ for some $T>0$, then $\dot{f}(T)=\zeta a(T)>0$ (since we know already that $a(T)>0$). But then, for $\varepsilon>0$ small enough, $f(T-\varepsilon)<f(T)=0$, which is a contradiction.
If instead $a(T)=1$, then $\dot{a}(T)=-\alpha f(T)-\beta b(T)\leqslant-\alpha f(T)<0$ (since we know already that $f(T)>0$). Hence, for $\varepsilon>0$ small enough, $a(T-\varepsilon)>a(T)=1$, which is absurd. This finishes the proof.
Appendix A
A.1. Proof of Lemma 3.1
Let $P(\lambda):=\lambda^{3}+B\lambda^{2}+C\lambda+D$. We claim that
(A.1) $P$ possesses two complex conjugate roots.
To check this, we point out that, since $P(\lambda)\geqslant D>0$ when $\lambda\geqslant 0$, we have that all the real roots of $P$ are negative.
Hence, if, for the sake of contradiction, the claim in (A.1) were false, then $P$ would have three real roots, say $-a$, $-b$ and $-c$, with $a$, $b$, $c\in(0,+\infty)$.
As a consequence,
$P(\lambda)=(\lambda+a)(\lambda+b)(\lambda+c)=\lambda^{3}+(a+b+c)\lambda^{2}+(ab+ac+bc)\lambda+abc$
and accordingly $B=a+b+c$, $C=ab+ac+bc$ and $D=abc$.
But then $BC=(a+b+c)(ab+ac+bc)>3D$, in contradiction with our assumption.
Having established (A.1), to complete the proof of Lemma 3.1 we need to show that when $BC<D$ the real part of these complex roots is positive.
To this end, from the above analysis we know that the three roots of $P$ have the form $-a$, $\alpha+i\beta$ and $\alpha-i\beta$, with $a\in(0,+\infty)$, $\alpha\in\mathbb{R}$, and $\beta\in\mathbb{R}\setminus\{0\}$, and our goal is to show that
(A.2) $\alpha\in(0,+\infty).$
For this, we observe that
$P(\lambda)$ $=$ $(\lambda+a)(\lambda-\alpha-i\beta)(\lambda-\alpha+i\beta)$
$=$ $\lambda^{3}+(a-2\alpha)\lambda^{2}+(\alpha^{2}-2a\alpha+\beta^{2})\lambda+a(\alpha^{2}+\beta^{2}).$
As a result,
$B=a-2\alpha,\quad C=\alpha^{2}-2a\alpha+\beta^{2}\quad{\mbox{and}}\quad D=a(\alpha^{2}+\beta^{2}).$
Consequently,
00 0 $<$ $D-BC$
$=$ $a(\alpha^{2}+\beta^{2})-(a-2\alpha)(\alpha^{2}-2a\alpha+\beta^{2})$
$=$ $2\alpha(a^{2}-2a\alpha+\alpha^{2}+\beta^{2})$
$=$ $2\alpha\big{(}(a-\alpha)^{2}+\beta^{2}\big{)},$
from which (A.2) plainly follows.
A.2. Proof of Lemma 3.2
We choose
(A.3) $f_{0}:=1,\qquad\gamma:=\eta:=c\alpha\qquad{\mbox{ and }}\qquad\zeta:=(1+c)\alpha,$
with $c\in(0,+\infty)$.
In this way, conditions (2.5) and (2.6) are satisfied and we see that
$B=(1+2c)\alpha,\qquad C=c(2+c)\alpha^{2}\qquad{\mbox{and}}\qquad E=-c\alpha\beta\theta,$
with $\beta\in(0,+\infty)$ and $\theta\in\mathbb{R}\setminus\{0\}$.
We now construct a case in which $\triangle>0$. For this, we suppose that $\alpha\to+\infty$ (for given $\beta$, $\theta$, and $c$) and we observe that (3.3) is fulfilled. In this situation, we take $c$ sufficiently small such that
$\Xi_{0}:=(1+2c)^{2}(2+c)^{2}-4c(2+c)^{3}>0.$
Then, we infer from (3.2) that
$\triangle=(1+2c)^{2}c^{2}(2+c)^{2}\alpha^{6}-4c^{3}(2+c)^{3}\alpha^{6}+O(\alpha^{4})=c^{2}\Xi\alpha^{6},$
where
$\Xi:=(1+2c)^{2}(2+c)^{2}-4c(2+c)^{3}+O\left(\frac{1}{\alpha^{2}}\right)=\Xi_{0}+O\left(\frac{1}{\alpha^{2}}\right).$
Hence, for $\alpha$ large enough (possibly depending on $\beta$, $\theta$, and $c$), we conclude that $\Xi\geqslant\frac{\Xi_{0}}{2}>0$ and accordingly $\triangle\geqslant\frac{c^{2}\Xi_{0}\alpha^{6}}{2}>0$.
Now we show that $\triangle<0$ is also possible. For this, we distinguish two cases, depending on the sign of $\theta$.
If $\theta\in(-\infty,0)$, we consider $\alpha:=1$, $\beta:=\frac{1}{|\theta|}$, and $c\in(0,+\infty)$ sufficiently small. In this situation, we note that
$\left(\alpha f_{0}+\gamma+\eta\right)\left(2\alpha f_{0}\eta+\gamma\eta\right)-\beta\eta|\theta|f_{0}=(1+2c)c(2+c)-c$
$\qquad=c\big{(}(1+2c)(2+c)-1\big{)}>c>0$
and therefore (3.3) is satisfied.
We also have that $B=1+2c$, $C=c(2+c)$ and $E=c$. Thus, we infer from (3.2) that
$\triangle=-4(1+2c)^{3}c+O(c^{2})<0,$
as long as $c$ is sufficiently small.
If instead $\theta\in(0,+\infty)$, then (3.3) is automatically satisfied. Thus, we consider $\alpha\searrow 0$ (for given $\beta$, $\theta$, and $c$) and we infer from (3.2) that
$\triangle=-27c^{2}\alpha^{2}\beta^{2}\theta^{2}+O(\alpha^{4})<0.$
These cases complete the proof of Lemma 3.2
Conclusions
We have introduced a model comprising a system of three ordinary differential equations to describe the interplay of bushfires events and prescribed burning activities in terms of available land.
By simple analytical methods, we have demonstrated that the model possesses a unique fixed point, representing a steady-state equilibrium of the system. The stability of this equilibrium is sensitive to the policy governing prescribed burning. Specifically:
  • Unstable Equilibrium with Incremental Prescribed Burning (Reactive Policy): When prescribed burning activities are increased in response to bushfire events, the equilibrium is unstable.
  • Stable Equilibrium with Prescribed Burning Replacing Natural Bushfires (Proactive Policy): In contrast, when moderate (not excessive) prescribed burning replaces natural bushfires, the equilibrium is stable and the system can maintain a balanced state. The stability of the equilibrium is also enhanced in the presence of a high regeneration factor.
We have also shown that the risks entailed in the Reactive Policy can be mitigated via a linear feedback control on the system.
Though our rigorous analytic results show how consistent policies are effective in bringing the system into a stable point, these results should not be seen as the definitive answer to the issue of prescribed burning. Many specific factors—such as climate conditions, vegetation types, fire history, and socio-economic implications—must be taken into account in any forward-thinking policymaking. However, the stability analysis presented here suggests that policies driven solely by reactions to bushfire events might be ineffective, potentially leading to instability. Therefore, despite their undeniable efficacy in raising public awareness, to ensure full effectiveness reactive policies should incorporate feedback control mechanisms. On the other hand, proactive policies that focus on planned and controlled prescribed burning, in moderate quantities and depending on recent bushfire intensity, could be more effective in prevention and long-term stability.